A theater group plans to sell 500 tickets for a play. They charge $45 for each orchestra seat and $30 for each balcony seat, and they plan to collect $18000 for a performance. How many of each type of tickets do they expect to sell?
Answer by mananth(16946)

x= number of orchestra seats
y= number of balcony seats

1 x + 1 y = 500 .............1
Total value
45 x + 30 y = 18,000 .............2
Eliminate y
multiply (1)by -30
Multiply (2) by 1
-30 x -30 y = -15,000
45 x + 30 y = 18,000
Add the two equations
15 x = 3,000
/ 15
x = 200
plug value of x in (1)
1 x + 1 y = 500
200 + y = 500
y = 500 -200
y = 300
y = 300
x= 200.00 number of orchestra seats
y= 300.00 number of balcony seats
m.ananth@hotmail.ca